Use Theorem 7.1.1 to find L{f(t)}. (Write your answer as a function of s.) f(t) = (2t - 1)3 %3D L{f(t)} =
Answer #1

Ans Given, f(t)= (2+-1) $(t) = 8-24464 - 1 : 1444) =เรีย (2446-18 =-38t? -เริเcts + 1548-L213 = 8+-12 เร} +6L48 - 2518 iu = (